OpenStudy (anonymous):

Integral of sin^3x/cosx dx

hartnn (hartnn):

split up \(\sin^3x dx= \sin^2x(\sin x dx)\) and plug in u = cos x dx du =... ?

OpenStudy (anonymous):

du = -sin x dx

hartnn (hartnn):

so now what does your integral change to ?

hartnn (hartnn):

use sin^2 = 1-cos^2 first

OpenStudy (anonymous):

\[\frac{ \sin^2x \sin x }{ \cos x } dx\] then \[\frac{ 1-\cos^2x \sin x}{ \cos x } dx ?\]

hartnn (hartnn):

don't forget to use the brackets \(\large \frac{( 1-\cos^2x )}{ \cos x }\sin x dx \)

hartnn (hartnn):

and now just plug in cos x = u sin x dx = -du

hartnn (hartnn):

did u get something like -(1-u^2)/u du can u integrate this ?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1-u^2 }{ u } du\]

hartnn (hartnn):

with the negative sign

OpenStudy (anonymous):

\[\int\limits_{}^{} u^{negative1} - u du\]

OpenStudy (anonymous):

with a negative sign

OpenStudy (anonymous):

\[-\ln u - \frac{ 1 }{ 2 }u^2\]

OpenStudy (anonymous):

So the answer: \[\ln cosx - \frac{ 1 }{ 2 }\cos^2x\]

hartnn (hartnn):

-(1/u - u) du -----> -ln u +u^2/2+c just sign errors

OpenStudy (anonymous):

So then \[-\ln cosx + \frac{ \cos^2x }{ 2 }\]

hartnn (hartnn):

+c yes, now correct :)

hartnn (hartnn):

- ln cos x can also be written as ln sec x ,i f you want to.

OpenStudy (anonymous):

Thank you! Does it need to be abs value?

hartnn (hartnn):

since cos can take negative values, but natural log (ln) cannot, we need to specify, the absolute value , like \(\large \ln|\cos x|\)

hartnn (hartnn):

or \(\large \ln |\sec x|\)

OpenStudy (anonymous):

Thanks again!!! =)

hartnn (hartnn):

welcome ^_^